During the constant pressure reversible non-flow process with air, the temperature increases from 400 0F to 1600 0F. Show process on the p-V and T-s coordinates and find the heat transfer in BTU/lbm.
1
Expert's answer
2020-05-19T09:13:26-0400

"T_1 = 400^o F"

"T_2 = 1600^o F"

The key point for solving this problem is : the heat transfer for a non-flow reversible constant pressure process equals to the change in enthalpy of the system. Therefore,

"q = c_p(T_2-T_1)"

regardless of the temperature, cp could be found by

"c_p=R.k\/k-1"

"R= 0.06856 BTU\/lb ^oF" The value for k also assumed to be 1.4.

Thus, "c_p = 0.2399= 0.24"

"=> Q=0.24(1600-400)= 288 BTU\/lb"